Show that the perpendiculars drawn from the origin to tangent planes to the cone π‘₯2 βˆ’ 𝑦2 + 5z2 + 4π‘₯𝑦 = 0 lie on the cone π‘₯2 βˆ’ 𝑦2 + 𝑧2 + 4π‘₯𝑦 = 0.Β 


1
Expert's answer
2022-02-10T13:06:07-0500

According to the definition of reciprocal cone, we want to find the reciprocal of

"\ud835\udc65^2 \u2212 \ud835\udc66^2 + 5z^2 + 4\ud835\udc65\ud835\udc66 = 0"

Equation of reciprocal of "\ud835\udc65^2 \u2212 \ud835\udc66^2 + 5z^2 + 4\ud835\udc65\ud835\udc66 = 0" is;

"Ax^2+By^2+Cz^2+2Fyz+2Gxz+2Hxy=0"

where,

"A=bc-f^2,\ B=ac-g^2,\ C=ab-h^2\\
F=gh-af,\ G=hf-bg,\ H=fg-ch"


Here, "a=1,\ b=-1,\ c=5,\ f=0,\ g=0,\ h=2".

So,

"A=bc-f^2=-5-0=-5,\ B=ac-g^2=5-0=5,\ C=ab-h^2=-1-4=-5\\
F=gh-af=0-0=0,\ G=hf-bg=0-0=0,\ H=fg-ch=0-10=-10"

Thus, equation of reciprocal of "\ud835\udc65^2 \u2212 \ud835\udc66^2 + 5z^2 + 4\ud835\udc65\ud835\udc66 = 0" is;

"-5x^2+5y^2-5z^2+(2\times0\times yz)+(2\times0\times xz)-20xy=0\\
\Rightarrow -5x^2+5y^2-5z^2-20xy=0\\"

Dividing through by -5;

"\displaystyle
\Rightarrow x^2-y^2+z^2+4xy=0,\ \text{which is the other cone given in the question.}"

Showing that that the perpendiculars drawn from the origin to tangent planes to the cone

"\ud835\udc65^2 \u2212 \ud835\udc66^2 + 5z^2 + 4\ud835\udc65\ud835\udc66 = 0" lie on the cone "\ud835\udc65^2 \u2212 \ud835\udc66^2 + \ud835\udc67^2 + 4\ud835\udc65\ud835\udc66 = 0" .